AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri Feb 4 12:49:54 2022
Author
*
:
Subject
*
:
JM plans to install new crimped, 3mm dia copper drain wires (a la AIDA@DESPEC) this afternoon. Initial Grounding - Lemo cables connecting adaptor cards ground in daisy chain. Fee bodies connected by 3mm copper drain wires in daisy chain. Two chains brought together at final adaptor card which is connected to cryring ground. Attachments 1-5 ADC data item rates all <10k ASIC settings all LEC slow comparators 0x64 1.8.L spectra aida05-08 widths 556, 511, ?, 570 ch FWHM respectively 1.8.W spectra 20us FSR Adaptor card grounding image 1.8L widths appear broader than previous day of ~300 width. Reason unknown, no change made to setup as yet. Second Grounding - 3mm copper drain wires connect each adaptor card to Fee body. Fee bodies daisy chained together using 3mm copper drain wire and connected to cryring ground. ADC data item rates all <10k ASIC settings all LEC slow comparators 0x64 Attachments 6-8 1.8.L spectra aida05-06 widths 620, 550 ch FWHM respectively, 1.1.L aida08 width 350 ch FWHM 1.8.W spectra 20us FSR. Some extra high frequency noise Adaptor card grounding Third Grounding - 3mm copper drain wires connect adaptor cards in daisy chains. Fee bodies daisy chained together using 3mm copper drain wire. Two chains brought together and connected to cryring ground. ADC data item rates all <10k ASIC settings all LEC slow comparators 0x64 1.8.L spectra aida05-08 similar to initial grounding ~500 ch FWHM Attachments 9-11 1.8.W spectra 20us FSR. AIDA5-8 channel rates. Rates dominated by noisy strips. Missing strips come and go as ASIC settings checked. Adaptor card grounding Some work still required to reduce noise in the system and reduce pulser widths. Pulser should also be checked on scope Gauge turned back ON. Bias 80V, 3.5 uA before. Bias 80V, 5.4 uA after. Bias turned off and on again. NO change to diode characteristic, current ramps high before coming down once bias on. Bias left ON 80V to see if stable when second gauge turned on and valves reopened. Pulser ON. Fee power OFF. Water cooling ON. No significant changes to waveforms, pulser width or rates with gauge ON. Other noise contributions probably outweigh effect of the gauge at this point. 3mm copper grounding cable and attachments, crimper and Yorkshire tea left in control room.
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Resubmit as new entry
Attachment 1:
Screenshot_from_2022-02-04_13-37-54.png
Original size: 1920x1014
Attachment 2:
Screenshot_from_2022-02-04_13-38-14.png
Original size: 1920x1014
Attachment 3:
Screenshot_from_2022-02-04_13-43-30.png
Original size: 1920x1014
Attachment 4:
Screenshot_from_2022-02-04_13-48-28.png
Original size: 1920x1014
Attachment 5:
20220203_100719.jpg
Original size: 4000x3000
Attachment 6:
2022-02-04_15-03-391.8*L.png
Attachment 7:
2022-02-04_14-57-50aida8-1.8*W.png
Attachment 8:
2022-02-04_16-13-39-1*W.png
Attachment 9:
2022-02-04_16-01-49rateAIDA5-8.png
Original size: 1920x1014
Attachment 10:
20220204_154113.jpg
Original size: 2250x3000
Attachment 11:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown